OpenStudy (anonymous):

please help: Use the substitution method to solve the linear system. 1/2x+y=2 2x+3y=9 different problem 1/3x+5/6y=1 1/2x-y=1

OpenStudy (precal):

@Desiree97 would you like me to show you how to do the first one?

OpenStudy (anonymous):

sure

OpenStudy (precal):

ok lets take (1/2)x+y=2 rewrite it to say y=? tell me what that would be

OpenStudy (anonymous):

y=1/2x+2

OpenStudy (precal):

when you move (1/2)x to the other side, you need to switch the sign to negative so y=2-(1/2)x

OpenStudy (precal):

Now sub this into the second equation

OpenStudy (precal):

what do you get?

OpenStudy (anonymous):

i have x-4=9 so x = 13

OpenStudy (precal):

no, 2x+3(2-(1/2)x)=9 solve for x tell me what you get

OpenStudy (precal):

waiting........

OpenStudy (anonymous):

x=5

OpenStudy (precal):

no 2x+3(2-(1/2)x)=9 distribut 3 2x+6-(3/2)x=9 -6 -6 (4/2)x-(3/2)x=3 (1/2)x=3 mulitply 2 to both sides x=6

OpenStudy (precal):

now that you know x=6 use 2x+3y=9 sub for x and tell me what y is

OpenStudy (precal):

waiting.....

OpenStudy (anonymous):

y=13

OpenStudy (precal):

2(6)+3y=9 solve for y....... your y value is incorrect

OpenStudy (anonymous):

y =7

OpenStudy (precal):

want to show some steps because once again you are incorrect

OpenStudy (anonymous):

yeah show me some steps

OpenStudy (precal):

no, I would like you to show me some steps first step is 2(6)+3y=9

OpenStudy (anonymous):